# Heterogeneous Treatment Effects and

### v2

@@ Abstract
[…]
The credibility revolution in empirical social science has motivated the largescale adoption of research designs that facilitate unbiased estimation of causal effects (Samii, 2016; Angrist and Pischke, 2010). Such unbiased estimates allow for valid measurement of the effect of a treatment on an outcome. However, they do not generally provide evidence about why or how the treatment affected the outcome. These questions of why or how are ultimately questions about the activation and influence of causal mechanisms. Understanding the mechanisms that produce causal effects is central to our ability to use empirical evidence to understand social phenomena (Slough and Tyson, 2024a).
Applied researchers typically pursue a number of different approaches to ascertaining the mechanisms that generate causal effects. There exist at least four distinct approaches in the applied literature: (1) evaluation of the sign of treatment effects on a given outcome (e.g., Ashworth, Berry, and de Mesquita, 2023); (2) mediation analysis (Imai, Keele, and Tingley, 2010; Imai et al., 2011; Glynn, 2012; Imai and Yamamoto, 2013); (3) multimethod research involving some form of qualitative or quantitative triangulation of causal findings (Levy Paluck, 2010; Dunning, 2012); and (4) the estimation of heterogeneous treatment effects (HTEs). The last approach—HTEs—compares estimated treatment effects for various subgroups thought to be informative about mechanism activation. While this is currently the modal approach to (quantitative) mechanism-testing in political science, the theoretical properties of this enterprise are not well explored.
To examine the prevalence of the use of HTEs to learn about causal mechanisms, we survey the 2021 volumes of three leading journals in political science: the American Journal of Political Science (AJPS), the American Political Science Review (APSR), and the Journal of Politics (JoP). We first identify the subset of papers that analyze quantitative empirical data. We then report the proportion of those empirical papers that report HTEs or subgroup-specific treatment effects. Finally, we calculate the proportion of papers using HTEs that interpret these quantities as providing information about causal mechanisms(s). Table 1 shows that in each of the three leading journals, a majority of quantitative studies estimate HTEs. Moreover, conditional on reporting any HTEs, the vast majority of articles (82% across three journals) interpret these quantities as providing information about mechanisms. Collectively, these figures indicate that almost half (46%) of recent quantitative empirical articles in these journals use HTEs to assess mechanisms. Table A.1 further documents that the share of studies that use of HTE for mechanism detection is similar across all quantitative research designs/identification strategies in common usage.
@@
Existing concerns about HTEs have largely focused on the statistical properties of relevant estimators and hypothesis tests. In particular, interaction terms in regression—the most common estimators of HTEs—are known to have low statistical power (e.g., McClelland and Judd, 1993). Moreover, estimation of HTEs with respect to many (pre-treatment) covariates risks multiple comparisons problems (Gerber and Green, 2012; Lee and Shaikh, 2014; Fink, McConnell, and Vollmer, 2014). While these criticisms are important, they are distinct from theoretical questions about how the presence or absence of HTEs links to causal mechanisms. We take on this challenge by asking: under what conditions do HTEs provide evidence of mechanism activation?
To answer this question, we develop a framework to formally link HTEs with respect to a covariate to the effect of a specific mechanism. To do so, we extend the workhorse causal mediation framework (Imai, Keele, and Tingley, 2010). Within our framework, a mechanism is an underlying process that influences experience in order to produce a (causal) effect if it is activated (Slough and Tyson, 2024a). A mediator, or mechanism representation, should thus be affected by treatment and have a non-zero (indirect) effect on the outcome if the mechanism is active for some unit. In order to use HTEs to detect mechanisms, empiricists rely on a measured moderator or pre-treatment covariate that is thought to predict the degree to which treatment activates a mechanism and/or the mechanism’s effect on an outcome of interest. Our results characterize the conditions under which heterogeneity in conditional average treatment effects (CATEs) with respect to a given moderator is a sufficient condition to show that the mechanism is active. A mechanism is active when its indirect effect is non-zero for some unit.
@@
We first present the exclusion assumption that links HTEs to a specified mechanism. The assumption holds that the moderator of interest is excluded with respect to other (average) effects of the treatment on the outcome. The intuition for the assumption is straightforward. The use of HTEs to assess the activation of a given mechanism relies on the additive separability of the effect of that mechanism from the effects of other mechanisms. The exclusion assumption is needed to ensure that additive separability is possible. In the absence of such an assumption, the relationship between HTEs and the indirect effect of a mechanism is unspecified. In this sense, this assumption is implicitly invoked by current practice, but is not explicitly stated or defended.
@@
When the exclusion assumption holds, what can we learn about a mechanism of interest from HTEs with respect to a given covariate? Our main results characterize what we can learn from the existence or non-existence of HTEs. We show that whether such learning is possible depends on the relationship between measured outcomes and the theory. We thus distinguish between outcome variables that are theorized to be directly affected by a mechanism versus (causally) subsequent outcomes that are indirectly affected.
First, consider the case in which a measured outcome is directly affected by the mechanism of interest. We show that when the exclusion assumption holds, the existence of HTEs provides evidence of mechanism activation. This broadly conforms to current practices in applied research, albeit while making the underlying assumption explicit. However, when there do not exist HTEs, there are two possible explanations: (1) the indirect effect of a mechanism is not moderated by any covariate; and/or (2) the theorized relationship between the covariate and the indirect effect of the mechanism was misspecified. Neither possibility distinguishes between an active or inactive mechanism. This means that a lack of HTEs cannot “rule out” a mechanism by showing that it is inactive. This finding contradicts standard interpretations of (the absence of) HTEs in the applied literature.
@@
We further consider the common case in which we observe outcomes that are indirectly affected by a mechanism. For example, in a political economy model, a mechanism may change an actor’s utility, but researchers measure only the actor’s discrete choice/behavior. In political psychology, a treatment may affect a subject’s (latent) attitudes, but researchers measure their survey response on a Likert scale. When an outcome is generated by a non-linear mapping of the directly-affected outcome (as in both of these examples), the relationship between HTEs and mechanisms changes. In this case, the existence of HTEs with respect to a covariate no longer provides evidence of mechanism activation, even when the exclusion assumption holds. The logic for this result is straightforward: the non-linear mapping of the mechanism’s influence into an observed outcome breaks the additive separability of the indirect effect of interest from other indirect and direct effects, which undermines our ability to link HTEs to the indirect effect of a mechanism.
This paper makes three principal contributions. First, our framework clarifies distinctions between varying uses of HTEs. Our focus, in line with most current empirical applications, is on the use of HTEs to learn about the mechanisms that produce treatment effects. This use of HTEs is distinct from important recent methodological contributions that emphasize the use of HTEs for extrapolation, prediction, or targeting of treatments (e.g., Egami and Hartman, 2022; Huang, 2022; Devaux and Egami, 2022; Kitagawa and Tetenov, 2018; Athey and Wager, 2021). Whereas these applications are facilitated by the application of machine learning methods (e.g., Grimmer, Messing, and Westwood, 2017; Chernozhukov et al., 2018; Athey, Tibshirani, and Wager, 2019), we show that using HTEs to learn about mechanisms instead relies on a deductive theoretical mapping of covariates to mechanisms, which is unlikely to be advanced by more widespread use of machine learning. Clarity about this distinction will improve interpretation of HTEs in applied work. It will further contribute to the development of methods for the multiple uses of HTEs.
@@
Second, we expand a growing literature on the theoretical implications of empirical models (TIEM) (Bueno de Mesquita and Tyson, 2020; Ashworth, Berry, and Bueno de Mesquita, 2021; Ashworth, Berry, and de Mesquita, 2023; Abramson, Koçak, and Magazinnik, 2022; Slough, 2023). We make two central interventions to this literature. First, our framework makes explicit links between a causal mediation framework that is used more prominently by empiricists and (formal) theoretical models. Second, we introduce questions about how measured outcomes relate to theoretical constructs by distinguishing between directly- and indirectly-affected outcomes. While measurement is central to recent TIEM work on evidence accumulation in a cross-study environment (Slough and Tyson, 2023, 2024a, 2024b), it has not been widely explored in the single-study environment.
@@ 1 Current Practice
Second, the absence of HTEs with respect to a given covariate is frequently used to “rule out” the activation of a possible mechanism (often called an alternative explanation). For example, Moscowitz (2021) argues that the proportion of in-state residents in one’s local media market increases rates of voter political knowledge and split-ticket voting in the US, countering trends toward the nationalization of politics. The paper provides evidence that news coverage is the mechanism that drives this effect. However, it also seeks to rule out an alternative mechanism: campaign advertising. The logic for this alternative mechanism holds that more in-state residents leads to more campaign advertising about in-state candidates, which in turn increases voter knowledge. Since advertisements air when during election season (when incumbents contest re-election), if the advertising mechanism were operative, we would expect the effect of in-state residents on voter knowledge to be larger during election seasons than other times in the electoral cycle. Moscowitz (2021) finds no such heterogeneity. He uses this finding is used to provide evidence against the advertisement-based mechanism.
Both of the above examples are exceptionally clear in delineating the mechanisms under consideration using HTEs, and thereby serve as exemplars of current practice. We show how current practice with respect to HTEs and the detection of causal mechanisms can mislead. We proceed from these empirical examples to a purely theoretical motivating example to illustrate the concern.
@@ 2 Motivating Example: Exogenous Shocks and Voting
In this formulation, higher values of $\omega$ correspond to a more adverse shock (e.g., a more intense natural disaster). Function $f$ is monotonically increasing in $\theta$ and decreasing in $\omega$. $\varepsilon$ is an idiosyncratic shock to the governance outcome that is drawn from a symmetric, continuously differentiable probability density function, $\phi$, that satisfies the monotone likelihood ratio property relative to $g$ [^4] Formally, this implies that if $x^{\prime}>x$, then $\frac{\phi(g-x^{\prime})}{\phi(g-x)}$ is strictly increasing in $g$.
@@
In the population, $v_{i}\sim U(-1,1)$. Voters have heterogeneous prior beliefs about the probability that incumbent is of type $\overline{\theta}$, formally $\pi_{i}(\overline{\theta})\in[0,1]$, where $\pi_{i}(\overline{\theta})\sim F_{\pi}$. For simplicity, we assume that voters share a common prior $\pi^{C}\in(0,1)$ for the challenger. [^5] The superscript $C$ denotes the challenger. Any belief without a superscript pertains to the incumbent. Voters vote either for the incumbent or challenger. The sequence is as follows:
@@ 2.1 From Theory to Empirical Research Design
Mapping this model onto the empirical research design, we will assume that $\omega\in\{\omega^{\prime},\omega^{\prime\prime}\}$ denotes a binary treatment, where $\omega^{\prime}$ indicates no exposure to the (adverse) shock and $\omega^{\prime\prime}$ denotes exposure to the shock. The first outcome, $y_{1}(\omega)$, measures voters’ expected utility from the incumbent. It is obviously difficult and rare to measure utility directly, though one could, in principle, elicit willingness-to-pay. A voter’s expected utility from a vote for the incumbent is given by:
@@
Our model can be represented as a directed acyclic graph (DAG), as depicted in Figure 1. This representation clarifies a number of assumptions of the model. First, the mechanism through which the shock affects voter preferences and behavior is voter learning. This is evident because the only path from the shock ($\omega$) to the outcomes $y_{1}(\omega)$ and $y_{2}(\omega)$ passes through the voters’ posterior beliefs about the incumbent’s type. Voter learning is moderated by voters’ prior beliefs about the incumbent, $\pi_{i}(\overline{\theta})$. Throughout this paper, we will indicate causal moderation through the arrow pointing to a path rather than a node, following the suggestion of Weinberg (2007). [^6] We could define an extra node to indicate this interaction. We do not do so in the interest of parsimony. Note that the representation of “interaction” effects in DAGs is not standardized (Nilsson et al., 2021). Because our theoretical model encodes more structure (e.g., excluding the possibility of certain causal moderation effects), this notational convention encodes more structure in graphical representations than is standard. For the purposes of exposition, the mediator, posterior beliefs, is unobserved to researchers. We will assume, however, that researchers have a measure of $\pi_{i}$ and $v_{i}$ from a baseline survey.
@@
Empirical researchers typically will not fully understand the causal structure represented in Figure 1. If this were the case, a researcher may mistakenly think that a shock affects assessments of valence, such that $v_{i}$ is a moderator. (For reference, we depict this alternative DAG that is inconsistent with our model in Figure A.1.) In the absence of measured mediators, the researcher could assess the mechanisms using by evaluating differences in conditional (or subgroup) treatment effects. Specifically, we will follow common practice by supposing that researchers estimate conditional average treatment effects (CATEs) at different levels of a moderator, $X$, as follows:
@@
Researchers will detect heterogeneity with moderator $\pi$ (by a). This occurs because the effect of the shock on a voter’s posterior belief is moderated by her prior in our model. Moreover, they will only detect heterogeneity in $\pi$, not in $v$, (by c) supporting the inference that the learning mechanism produces the observed ATE. In contrast, HTEs are not observed for the (non)-moderator $v$ (by b) because valence and posterior beliefs (the mechanism) are additively separable in voters’ utility. Here, researchers are unlikely to mis-attribute the mechanism through analysis of HTEs.

### v3

@@ Abstract
Social scientists often make claims about how or why a treatment affects a given outcome. Researchers might tell us, for example, not only that an outreach program increased COVID-19 reporting rates, but that it did so by building trust in government. Or researchers might claim that access to local media increases knowledge of local politics because of the content of news coverage, not because of local political advertising. Claims of this type are often made on the basis of heterogeneous treatment effects (HTEs): the effect of the outreach program on COVID-19 reporting is larger in places that initially distrusted government (Haim, Ravanilla, and Sexton, 2021), for example, or the effect of local media on knowledge of local politics does not change over the election cycle (Moscowitz, 2021). In fact, a majority of recent articles that report HTEs interpret them as evidence about how or why a treatment affects a given outcome.
But the theoretical foundation for making such claims lags far behind empirical practice. While we have strong theoretical foundations for making claims about treatment effects themselves, theory provides little guidance about the relationship between HTEs and tests of mechanisms. We provide that guidance, developing a theoretical framework that clarifies the assumptions that researchers need to make in order to use HTEs as evidence for or against potential mechanisms. Our analysis reveals that current practice is often misleading—but also that, as is so often the case, making implicit assumptions explicit can correct such errors in future work.
This framework is important because political scientists so often use HTEs to make claims about causal mechanisms. Surveying the 2021 volumes of three leading journals in political science—the American Journal of Political Science (AJPS), the American Political Science Review (APSR), and the Journal of Politics (JoP)—–we find that a majority (56%) of quantitative studies estimate HTEs and that, conditional on reporting any HTEs, the vast majority of articles (82%) interpret them as providing information about mechanisms. Taken together, these figures indicate that almost half (46%) of recent quantitative empirical articles in these journals use HTEs to assess mechanisms. [^3] Table A.1 further documents that the share of studies that use of HTE for mechanism detection is similar across all quantitative research designs/identification strategies in common usage. Beyond our analysis, Blackwell, Ma, and Opacic (2024: p. 3) show that HTEs are the modal way that authors in these journals analyze causal mechanisms.
@@
We ask an essential but as-yet-unanswered question: under what conditions do HTEs provide evidence of mechanism activation? We do so by extending the workhorse causal mediation framework (Imai, Keele, and Tingley, 2010). We define a mechanism as an underlying process that influences experience in order to produce a (causal) effect if it is activated (Slough and Tyson, 2024a). In the context of a running example about the effect of adverse shocks like a natural disaster on pro-incumbent voting, voter learning about an incumbent has been posited as a potential mechanism by which such shocks affect voters’ decisions at the ballot box (e.g., Ashworth, Bueno de Mesquita, and Friedenberg, 2018). Whereas the mechanism itself is an unobserved theoretical construct, a mediator is its empirical counterpart: the voter’s posterior belief about the incumbent’s compentence, for example. If the learning mechanism is active, the mediator (a posterior belief) should be affected by the treatment (the negative shock), and it should have a non-zero (indirect) effect on the outcome (vote choice). In order to use HTEs to detect mechanisms, empiricists rely on a measured moderator, or pre-treatment covariate, that is thought to predict the degree to which treatment activates a mechanism and/or the degree to which the mechanism affects on the outcome of interest. In our running example, voters’ prior beliefs about the incumbent’s competence could be a moderator: prior beliefs should condition the extent of learning from the negative shock. Researchers might therefore argue that the existence of HTEs of the adverse shock on pro-incumbent voting at different levels of prior beliefs (the moderator) provide evidence in favor of the voter learning mechanism.
Our results characterize the conditions under which such conclusions are valid, i.e., the conditions under which heterogeneity in conditional average treatment effects (CATEs) with respect to a given moderator is sufficient to show that the mechanism is active. We consider a mechanism to be active when its indirect effect is non-zero for some unit.)
@@
The first such condition is that the moderator of interest is excluded with respect to other (average) effects of the treatment on the outcome. If, for example, voter prior beliefs (the moderator) were to condition the effect of an adverse shock on pro-incumbent voting via mechanisms other than voter learning—such as via directional motives or motivated reasoning (Little, Schnakenberg, and Turner, 2022)—then the exclusion assumption would be violated. The intuition is straightforward: The use of HTEs to assess the activation of a given mechanism relies on the additive separability of the effect of that mechanism from the effects of other mechanisms. In the absence of this exclusion assumption, the relationship between HTEs and the indirect effect of a mechanism is unspecified. This assumption is implicitly invoked by current practice but is rarely explicitly stated or defended.
@@
The second condition concerns the relationship between measured outcomes and the theory. When the exclusion assumption does hold, what can we infer about a mechanism of interest from HTEs with respect to a given moderator (covariate)? Our main results characterize what we can learn from the presence or absence of HTEs. We show that whether such inferences are justified depends on the relationship between measured outcomes and the theory. We thus distinguish between outcome variables that are theorized to be directly affected by a mechanism versus (causally) subsequent outcomes that are indirectly affected.
First, consider the case in which a measured outcome is directly affected by the mechanism of interest. This would be the case if, for example, a researcher was measuring voter utility using an elicitation of voter willingness to pay for the incumbent over the challenger. We show that when the exclusion assumption holds, the existence of HTEs does provide evidence of mechanism activation. This result broadly supports current practices in applied research, albeit while making the underlying assumption explicit. But the absence of HTEs does not rule out mechanism activation. Indeed, the absence of HTEs could arise because: (1) the mechanism is inert (e.g., voters do not learn about the incumbent’s competence from a negative shock); (2) the moderator does not in fact affect the mechanism (e.g., a voter’s prior belief does not condition how much she learns from a shock); and/or (3) because the mechanism produces the same effect on all units (e.g., voter learning produces the same effect on all voters’ assessment of the incumbent). We cannot distinguish between these three possibilities from a lack of HTEs, meaning that we cannot use the absence of HTEs cannot “rule out” a mechanism. This finding invalidates common interpretations of the absence of HTEs in the applied literature.
@@
We further consider the common case in which researchers observe outcomes that are indirectly affected by a mechanism. For example, as is standard in studies of exogenous shocks and voting, researchers measure vote choice. In these cases, a treatment may affect a voter’s (latent) utility, but researchers measure only their choice of incumbent or challenger. When there is a non-linear mapping from the directly affected outcome to the observed, indirectly affected outcome (as in this example), the relationship between HTEs and mechanisms changes. In this case, the existence of HTEs with respect to a moderator no longer provides evidence of mechanism activation, even when the exclusion assumption holds. The logic for this result is straightforward: the non-linear mapping breaks the additive separability of the indirect effect of interest from other indirect and direct effects, which undermines our ability to link HTEs to the indirect effect of a mechanism. This means that even if voters respond to a negative shock only via motivated reasoning—not Bayesian learning—there will generally exist HTE on vote choice when using voters’ prior beliefs as a moderator. This means that the existence of HTEs does not provide evidence of voter learning as a mechanism. Again, this finding counters common interpretations of the presence of HTE for indirectly-affected outcomes.
This paper makes three principal contributions. First, we introduce a new framework to understand the theoretical relationship between causal mechanisms and treatment effect heterogeneity. A large methodological literature provides guidance on the estimation of hetergeneous treatment effects (or “interaction effects”) (e.g., Brambor, Clark, and Golder, 2017; Berry, DeMeritt, and Esarey, 2009; Hainmueller, Mummolo, and Xu, 2018; Grimmer, Messing, and Westwood, 2017; Athey, Tibshirani, and Wager, 2019). More recent contributions the use of HTEs for extrapolation, prediction, or targeting of treatments (e.g., Egami and Hartman, 2022; Huang, 2022; Devaux and Egami, 2022; Kitagawa and Tetenov, 2018; Athey and Wager, 2021). Yet, because these contributions are primarily statistical, they do not facilitate theoretical analysis of the relationship between mechanisms and the (measured) treatment effects that they produce. Our primary results on the theoretical properties of HTEs—and the problems with existing practice—therefore complement our understanding important statistical issues associated with the estimation of HTEs, namely limited statistical power (e.g., McClelland and Judd, 1993) and multiple-comparisons problems (Gerber and Green, 2012; Lee and Shaikh, 2014; Fink, McConnell, and Vollmer, 2014).
@@
Second, we expand a growing literature on the theoretical implications of empirical models (TIEM) (Bueno de Mesquita and Tyson, 2020; Ashworth, Berry, and Bueno de Mesquita, 2021; Ashworth, Berry, and de Mesquita, 2023; Abramson, Koçak, and Magazinnik, 2022; Slough, 2023). We make two central interventions to this literature. First, our framework makes explicit links between a causal mediation framework that is used more prominently by empiricists and (formal) theoretical models. This analysis complements recent work by Blackwell, Ma, and Opacic (2024), who make explicit the strong assumptions that underpin the use of treatment effects on intermediate outcomes to discern causal mechanisms. Second, we introduce questions about how measured outcomes relate to theoretical constructs by distinguishing between directly- and indirectly-affected outcomes. While measurement is central to recent TIEM work on evidence accumulation in a cross-study environment (Slough and Tyson, 2023, 2024a, 2024b), it has not been widely explored in the single-study environment.
@@ 1 Current Practice
Second, the absence of HTEs with respect to a given covariate is frequently used to “rule out” the activation of a possible mechanism (often called an alternative explanation). For example, Moscowitz (2021) argues that the proportion of in-state residents in one’s local media market increases rates of voter political knowledge and split-ticket voting in the US, countering trends toward the nationalization of politics. The paper provides evidence that news coverage is the mechanism that drives this effect. However, it also seeks to rule out an alternative mechanism: campaign advertising. The logic for this alternative mechanism holds that more in-state residents leads to more campaign advertising about in-state candidates, which in turn increases voter knowledge. Since advertisements air when during election season (when incumbents contest re-election), if the advertising mechanism were operative, we would expect the effect of in-state residents on voter knowledge to be larger during election seasons than other times in the electoral cycle. Moscowitz (2021) finds no such heterogeneity. The absence of heterogeneity is used to provide evidence against the advertisement-based mechanism. Both of the above examples are exceptionally clear in delineating the mechanisms under consideration using HTEs, and therefore serve as exemplars of current practice.
@@ 2 Motivating Example: Exogenous Shocks and Voting
In this formulation, higher values of $\omega$ correspond to a more adverse shock (e.g., a more intense natural disaster). Function $f$ is monotonically increasing in $\theta$ and decreasing in $\omega$. $\varepsilon$ is an idiosyncratic shock to the governance outcome that is drawn from a symmetric, continuously differentiable probability density function, $\phi$, that satisfies the monotone likelihood ratio property relative to $g$ [^5] This implies that if $x^{\prime}>x$, then $\frac{\phi(g-x^{\prime})}{\phi(g-x)}$ is strictly increasing in $g$.
@@
In the population, $v_{i}\sim U(-1,1)$. Voters have heterogeneous prior beliefs about the probability that incumbent is of type $\overline{\theta}$, formally $\pi_{i}(\overline{\theta})\in[0,1]$, where $\pi_{i}(\overline{\theta})\sim F_{\pi}$. For simplicity, voters share a common prior $\pi^{C}\in(0,1)$ for the challenger. [^6] The superscript $C$ denotes the challenger. Any belief without a superscript pertains to the incumbent. This simplifying assumption of asymmetry in voters’ priors about the incumbent versus challenger allows us to focus on core related to interpretation of HTEs. However, such a scenario could plausibly be generated by varying attention to the news or governance outcomes during the incumbent’s term (prior to the shock). Voters vote either for the incumbent or challenger. The sequence is as follows:
@@ 2.1 From Theory to Empirical Research Design
Mapping this model onto the empirical research design, we will assume that $\omega\in\{\omega^{\prime},\omega^{\prime\prime}\}$ denotes a binary treatment, where $\omega^{\prime}$ indicates no exposure to the (adverse) shock and $\omega^{\prime\prime}$ denotes exposure to the shock. The first outcome, $y_{1}(\omega)$, measures voters’ expected utility from the incumbent. It is obviously difficult and rare to measure utility directly, though one could, in principle, elicit willingness-to-pay for the incumbent over the challenger. A voter’s expected utility from a vote for the incumbent is given by:
@@
Our model can be represented as a directed acyclic graph (DAG), as depicted in Figure 1. This representation clarifies a number of assumptions of the model. First, the mechanism through which the shock affects voter preferences and behavior is voter learning. This is evident because the only path from the shock ($\omega$) to the outcomes $y_{1}(\omega)$ and $y_{2}(\omega)$ passes through the voters’ posterior beliefs about the incumbent’s type. Voter learning is moderated by voters’ prior beliefs about the incumbent, $\pi_{i}(\overline{\theta})$. Throughout this paper, we will indicate causal moderation through the arrow pointing to a path rather than a node, following the suggestion of Weinberg (2007). [^7] We could define an extra node to indicate this interaction. We do not do so in the interest of parsimony. Our choice reflects the fact that representation of “interaction” effects in DAGs is not standardized (Nilsson et al., 2021). Because our theoretical model encodes more structure (e.g., excluding the possibility of certain causal moderation effects), this notational convention encodes more structure in graphical representations than is standard. For the purposes of exposition, the mediator, posterior beliefs, is unobserved to researchers. This feature is standard in most applied research on adverse shocks and voting behavior and renders explicit mediation analysis infeasible. We will assume, however, that researchers have a measure of $\pi_{i}$ and $v_{i}$ from a baseline survey.
@@
Empirical researchers typically will not fully understand the causal structure represented in Figure 1. If this were the case, a researcher may mistakenly think that a shock affects assessments of valence, such that the baseline measure of $v_{i}$ is a moderator. (For reference, we depict this alternative DAG that is inconsistent with our model in Figure A.1.) In the absence of measured mediators, the researcher could assess the mechanisms by evaluating differences in conditional (or subgroup) treatment effects. Specifically, we will follow common practice by supposing that researchers estimate conditional average treatment effects (CATEs) at different levels of a moderator, $X$, as follows:
@@
Researchers will detect heterogeneity with moderator $\pi$ (by a). This occurs because the effect of the shock on a voter’s posterior belief is moderated by her prior in our model. This is a straightforwad implication of Bayesian updating. Moreover, they will only detect heterogeneity in $\pi$, not in $v$, (by c) supporting the inference that the learning mechanism produces the observed ATE. In contrast, HTEs are not observed for the (non)-moderator $v$ (by b) because valence and posterior beliefs (the mechanism) are additively separable in voters’ utility. Here, researchers are unlikely to mis-attribute the mechanism through analysis of HTEs.